# TerraCodec: Compressing Earth Observations

### v1

@@ 1 Introduction
The exponential growth of Earth Observation (EO) data, driven by initiatives such as the Copernicus program, creates critical bottlenecks in storage, transmission, and processing (Guo et al., 2016; Wilkinson et al., 2024). EO imagery also differs fundamentally from natural images. It is multispectral, with up to dozens of channels beyond the visible RGB spectrum; and multi-temporal, with images captured at regular intervals from near-constant viewpoints but subject to seasonal, atmospheric, and cloud cover changes. As a result, EO scenes contain strong spatial and spectral redundancy, while temporal evolution arises from recurring seasonal patterns rather than object motion. These properties create compression challenges distinct from natural images, but make EO well-suited for learned approaches that capture domain-specific priors (Gomes et al., 2025). Despite advances in neural codecs for natural images and video (Ballé et al., 2017; Agustsson et al., 2020), such methods are not aligned with EO. Image codecs focus on RGB inputs and overlook temporal redundancy, while video codecs often rely on motion priors that are less suited for static landscapes with low temporal resolution.
@@
We address these gaps with TerraCodec, a family of neural codecs tailored to EO data. TerraCodec includes efficient image-based models for multispectral inputs: a lightweight Factorized Prior variant (TEC-FP) and an ELIC-based variant (TEC-ELIC) for optimal rate-distortion. We further introduce a Temporal Transformer (TEC-TT) that captures long-range dependencies across time without relying on hand-crafted motion priors. As shown in Figure 1, these models offer a significantly better reconstruction than standard codecs at the same compression rate. At equal image quality, they reduce storage size by up to an order of magnitude, with TEC-TT providing further reduction on long time series by exploiting temporal structure.
Our main contributions are: (1) TerraCodec, a suite of learned multispectral and multi-temporal codecs that achieve superior rate–distortion on EO data; (2) Latent Repacking, a method to train variable-rate neural codecs which we demonstrate with our FlexTEC model; and (3) downstream evaluations, demonstrating the utility of compression models for downstream tasks and zero-shot cloud inpainting. We release code and pretrained weights under a permissive license to support future research and adoption.
@@ 2 Related work
Foundations. Shannon’s source coding theorem bounds lossless compression by the source entropy; practical schemes such as Huffman and arithmetic coding approach this limit (Shannon, 1948; Huffman, 1952; Rissanen & Langdon, 1979). Lossy compression, in contrast, reduces storage requirements by discarding information. The rate–distortion function characterizes the minimum bitrate for a given distortion, formalizing the trade-off between rate and fidelity (Shannon, 1948). These principles underpin transform coding, which applies DCT or wavelets prior to quantization and entropy coding, forming the basis of standards like JPEG, JPEG2000, and HEVC (x265) (Ahmed et al., 1974; Daubechies, 1992; Wallace, 1991; Taubman & Marcellin, 2002; Sullivan et al., 2012).
@@
Neural compression. Learned codecs replace hand-crafted transforms with autoencoders trained end-to-end under a rate–distortion loss (Ballé et al., 2017; Theis et al., 2017). Inputs are mapped to latents, quantized, and entropy-coded under a learned prior. While recent work explores stronger backbones from convolutional models to transformers and diffusion-based decoders (Zhu et al., 2022; Li et al., 2024a; Yang & Mandt, 2023) and richer objectives (perceptual, adversarial) (Blau & Michaeli, 2019; Mentzer et al., 2020), a key performance trade-off is governed by the entropy model. Fully factorized priors offer efficiency (Ballé et al., 2017); hyperpriors (Ballé et al., 2018) introduce side information to capture spatially varying scales; autoregressive priors (Minnen et al., 2018) exploit local context at the cost of sequential decoding; and models such as ELIC (Cheng et al., 2020; He et al., 2022) additionally utilize efficient, parallel space–channel context. However, these image codecs are limited to a single rate–distortion setting per checkpoint. In contrast, flexible-rate models are using approaches such as conditioning on the rate parameter (Choi et al., 2019), spatially adaptive quality maps (Song et al., 2021; Tong et al., 2023), and hierarchical VAEs with quantization-aware priors (Duan et al., 2023).
@@
Beyond images, learned video compression targets temporal redundancy across frames (Agustsson et al., 2020; Li et al., 2023; 2024b). While classical approaches rely on motion estimation and compensation, transformer-based models remove such priors and model temporal dependencies directly in latent space. The Video Compression Transformer (VCT) (Mentzer et al., 2022) follows this design, encoding frames independently and using a temporal transformer to predict latents from past context, making it better suited to settings with limited or irregular motion.
@@
Earth Observation data. Most neural compression targets natural imagery, whereas EO includes multispectral bands, higher bit depth, and long temporal horizons. Compression must preserve spectral and structural cues relevant for downstream analysis (Gomes et al., 2025), aligning with the broader paradigm of task-oriented compression (Torfason et al., 2018; Singh et al., 2020). Operational EO pipelines typically rely on the JPEG2000 and CCSDS standards (Yeh et al., 2005; CCS, 2012; 2017) for their robustness and low complexity. Learned models have been explored for optical and SAR (Maharjan & Li, 2023; Di et al., 2022) images, with a focus on reducing on-board complexity (Alves de Oliveira et al., 2021) and spectral grouping (i Verdú et al., 2023). Other works exploit spatial–spectral encoders and mixed hyperpriors to capture redundancy (Kong et al., 2021b; a; Cao et al., 2022; Xiang & Liang, 2023; Fu & Du, 2023; Gao et al., 2023). Despite recent progress, learned EO compression remains fragmented, with most works targeting single images and providing little to no temporal modeling. To our knowledge, no pretrained models are publicly available for the widely used Sentinel–2 imagery. In contrast, TerraCodec offers efficient multispectral image codecs, a temporal transformer to capture long-range dependencies, and single-checkpoint, flexible-rate compression.
@@ 3 Methodology
We begin with an overview of our EO compression approach, then detail the architectures of the TerraCodec models 3.1, and finally introduce Latent Repacking for flexible-rate models 3.2.
We study lossy compression of multispectral, multi-temporal EO imagery. An EO sequence is a set of images $\mathbf{x}_{i}\in\mathbb{R}^{H\times W\times C}$, each of size $H\times W$ with $C$ spectral bands. While EO sensors range from a single panchromatic channel to finely sliced hyperspectral imagers, we focus on Sentinel–2 L2A with C=12C{=}12 optical bands in the visible and near infrared, saved as 16-bit radiometry.
@@
EO-specific choices. Our codecs adopt three EO-specific design choices: (i) native support for 12-band, 16-bit inputs; (ii) pretraining on a large-scale global EO dataset; and (iii) per-band standardization rather than global normalization, to stabilize training and preserve band-specific statistics for downstream tasks. Following prior literature (Ballé et al., 2017; Mentzer et al., 2022), we adopt CNN-based encoders for intra-frame compression due to their efficiency and low latency.
@@ 3.1 TerraCodec
We introduce TerraCodec, a family of learned codecs for EO, including two image codecs: a lightweight factorized prior (TEC-FP), a stronger space–channel context model (TEC-ELIC), and a temporal transformer (TEC-TT), with a flexible-rate variant (FlexTEC).
Figure 2: TerraCodec image codecs. Factorized Prior (Ballé et al., 2017) uses a fully factorized prior without context, and assumes a zero-centered normal distribution for the latents. ELIC (He et al., 2022) augments a hyperprior with spatial and channel context to predict per-latent mean/scale.
Factorized Prior (TEC-FP). TEC-FP is a Factorized Prior model (Ballé et al., 2017), our most basic TEC image codec. It employs a fully factorized entropy model, where each element of the quantized latent $\hat{\mathbf{y}}$ is modeled independently by $q_{\phi}(\hat{y}_{j})$, without side information or context. It is illustrated in the upper part of Figure 2. This yields fast, parallel entropy coding, but limited expressiveness compared to hyperprior- or context-based models.
Efficient Learned Image Compression (TEC-ELIC). TEC-ELIC instantiates ELIC’s space–channel context entropy model (He et al., 2022) for EO inputs. The encoder/decoder networks include residual bottleneck and attention blocks, increasing representational capacity. The entropy model predicts per-latent mean and scale from (i) spatial context via checkerboard convolutions, (ii) channel context from previously decoded latent groups, and (iii) side information from a hyperprior, improving rate–distortion performance at the cost of higher complexity. Figure 2 illustrates the hyperprior-based context model in simplified form.
Figure 3: Architecture of the TerraCodec-$T$ model following Mentzer et al. (2022). Each image is encoded with ELIC followed by a temporal transformer that predicts mean and scale for new images.
Temporal Transformer (TEC-TT). TEC-TT builds on the VCT architecture (Mentzer et al., 2022). We train a transformer to model temporal dependencies of seasonal EO data in latent space, predicting the current frame’s latent distribution from past context. Each frame $\mathbf{x}_{i}$ is encoded to latents $\mathbf{y}_{i}$ and quantized. We partition the current latent into $B$ non-overlapping spatial blocks $\{\hat{\mathbf{y}}_{i,b}\}_{b=1}^{B}$ and the two past latents into overlapping context blocks to increase the the transformer’s receptive field. Each block $b$ is flattened into a sequence of $T$ tokens $\{\hat{y}_{i,b,t}\}_{t=1}^{T}$ of channel width $d_{\text{lat}}$. A temporal encoder aggregates the two previous frames into a joint context embedding $z_{\mathrm{joint}}=E(\hat{\mathbf{y}}_{i-2},\hat{\mathbf{y}}_{i-1})$. As shown in Figure 3, within each current block, a masked autoregressive transformer predicts token-wise prior parameters conditioned on already decoded tokens and $z_{\mathrm{joint}}$ following Eq. 1.
@@
We assume conditional independence across blocks given the context, allowing parallel probability estimation during encoding and parallel block decoding. Causal masking prevents attention to undecoded tokens; see Appendix B.3 for details. TEC-TT uses the same CNN analysis–synthesis transforms as TEC-ELIC. Unlike the original VCT, it is trained end-to-end on the rate–distortion objective without image pretraining, using a $\lambda$-schedule that emphasizes low-rate regimes early. We further adapt TEC-TT for flexible-rate scaling, introducing the FlexTEC variant in the next section.
@@ 3.2 Latent Repacking for flexible-rate models
From spatial tokens to channel slices. A standard transformer codec represents an image block with $T$ spatial tokens, each spanning the full latent dimension $d_{\text{lat}}$. Dropping tokens at inference discards entire regions, causing severe artifacts (see App. F.2.1). Instead, we aim for early tokens to encode information that is globally useful across the scene.
Think of the latent block as a 3D tensor $A\in\mathbb{R}^{H\times W\times d_{\text{lat}}},$ with $H\times W=T$ tokens and latent dimension $d_{\text{lat}}$. In the standard layout, each token $t$ is a spatial patch $(h,w)$ containing all $d_{\text{lat}}$ channels at that location. Latent repacking instead slices the channel axis into $T$ groups of width $k=\tfrac{d_{\text{lat}}}{T},$ and redefines tokens so that each spans the full scene but only $k$ channels.
@@
In other words, $t^{\prime}_{u}$ is the $u$-th slice of $k$ consecutive channels of $A$, spanning the full spatial field. The procedure is reversible; reapplying the slicing and repacking restores the original layout. After repacking, keeping the first $K$ tokens corresponds to $A[:,:,0:K\cdot k],$ i.e. the first $K\cdot k$ latent channels at every spatial location.
@@
To keep later tokens informative, we use dynamic rate scaling: when fewer tokens are kept, their rate loss weight is upweighted. It prevents information from collapsing into the first tokens and encourages useful content across all tokens. Budgets $K$ are sampled more frequently at higher values (as in Bachmann et al. (2025)), ensuring that all tokens are trained while the model also learns to operate across a range of rates.
We apply the approach on TEC-TT and introduce the Flexible-Rate TerraCodec (FlexTEC) model. The model applies Latent Repacking and masking inside the temporal transformer after image-wise compression and restores the original layout before image decoding (details in App. C).
@@ 4 Experimental setup
This section describes the data and pretraining (4.1), evaluation and baselines (4.2), and downstream tasks (4.3) used to train and assess TerraCodec.
@@ 4.1 Pretraining
All TEC models are pretrained on SSL4EO-S12 v1.1 (Blumenstiel et al., 2025; Wang et al., 2023), a large-scale Sentinel–2 corpus with 244k globally distributed locations and four seasonal snapshots per location. Each L2A sample consists of $\times264$ pixels at 10 m resolution; we crop to $\times256$ pixels (random crops for training, center crops for evaluation) to ensure uniform size. Bands at 20 m and 60 m are upsampled to 10 m with nearest-neighbor interpolation for spatial alignment. We follow the official spatial split into training and validation sets. To stabilize multispectral training, each input band $b$ is standardized by its dataset mean $\mu_{b}$ and standard deviation $\sigma_{b}$. Losses are computed in this standardized space, which balances gradient magnitudes across channels and avoids overfitting to high-variance bands.
Image codecs. TEC-FP and TEC-ELIC are optimized with Adam ($\text{lr}=10^{-4}$), using an auxiliary learning rate of $5\cdot 10^{-3}$ for the entropy bottleneck, gradient clipping at 1.0, and mixed precision. We employ a cosine learning-rate schedule with 5% warmup and $\eta_{\min}=10^{-5}$. Models are trained for 100 epochs with batch size 64 on a single NVIDIA A100 GPU, requiring 20–25 hours. A temporal index is randomly sampled for each sample in every epoch, so that one epoch covers one quarter of the dataset. We sweep five $\lambda$ values to span low- to high-bitrate regimes.
Temporal codec. TEC-TT is trained with a temporal context of two past frames for 300k steps with a global batch size of 24 on four NVIDIA A100 GPUs, requiring about 70 hours. We optimize with AdamW ($\text{lr}\!=\!10^{-4}$, weight decay $10^{-2}$) and employ half-cosine decay schedule to $\eta_{\min}=10^{-6}$ with 15% warmup steps. We sweep six $\lambda$ values for varying rate–distortion settings. For low-rate settings ($\lambda\leq 5.0$), we scale $\lambda$ by 10 during the first 15% of training before annealing back to the target value. Further pretraining details, including FlexTEC, are given in Appendix B.

### v2

# TerraCodec: Compressing Optical Earth Observation Data

@@ 1 Introduction
The exponential growth of Earth observation (EO) data, driven by initiatives such as the Copernicus program, creates critical bottlenecks in storage, transmission, and processing [25, 57]. EO imagery also differs fundamentally from natural images. It is multispectral, with up to dozens of channels beyond the visible RGB spectrum; and multi-temporal, with images captured at regular intervals from nearly constant viewpoints. As a result, EO scenes contain strong spatial and spectral redundancy, while temporal evolution arises from recurring seasonal patterns rather than object motion. These properties create compression challenges distinct from natural images but make EO well-suited for learned approaches that capture domain-specific priors [23]. Despite advances in neural codecs for natural images and video [5, 1], neural compression for EO remains fragmented, with no available large-scale pretrained codecs for multispectral imagery, and temporal redundancy in satellite time series still largely unexplored.
@@
We address these gaps with TerraCodec, a family of neural codecs tailored to multispectral EO and pretrained on Sentinel-2 data. TerraCodec includes efficient image-based models for multispectral inputs: a lightweight Factorized Prior variant (TEC-FP) and an ELIC-based variant (TEC-ELIC) for optimal rate-distortion. We further introduce a Temporal Transformer (TEC-TT) that captures long-range dependencies across time without relying on hand-crafted motion priors. As shown in Fig. 1, these models offer a significantly better reconstruction than standard codecs at the same compression rate. At equal image quality, they reduce storage size by up to an order of magnitude, TEC-TT provides a further reduction for long time series by exploiting temporal structure.
Our main contributions are: (1) TerraCodec, a suite of Sentinel-2 pretrained multispectral and multi-temporal codecs that achieve superior rate–distortion performance over classical codecs; (2) Latent Repacking, a method to train variable-rate neural codecs which we demonstrate with our FlexTEC model; and (3) downstream evaluations, demonstrating the utility of compression models for analysis and zero-shot cloud inpainting. We release code and pretrained weights under a permissive license to support future research and adoption.
@@ 2 Related work
Foundations. Shannon’s source coding theorem bounds lossless compression by the source entropy; practical schemes such as Huffman and arithmetic coding approach this limit [45, 27, 44]. Lossy compression, in contrast, reduces storage requirements by discarding information. The rate–distortion function characterizes the minimum bitrate for a given distortion, formalizing the trade-off between rate and fidelity [45]. These principles underpin transform coding, which applies DCT or wavelets prior to quantization and entropy coding, forming the basis of standards like JPEG, JPEG2000, and HEVC (x265) [2, 17, 55, 50, 48].
@@
Neural compression. Learned codecs replace hand-crafted transforms with autoencoders trained end-to-end under a rate–distortion loss [5, 51]. Inputs are mapped to latents, quantized, and entropy-coded under a learned prior. Recent work extends this beyond convolutional autoencoders, exploring transformer backbones [67, 32], generative decoders [59, 36], and latent diffusion models [65], as well as richer perceptual and adversarial objectives [8, 40]. However, a key performance trade-off is governed by the entropy model. Fully factorized priors offer efficiency [5]; hyperpriors [6] introduce side information to capture spatially varying scales; autoregressive priors [41] exploit local context at the cost of sequential decoding; and models such as ELIC [14, 26] additionally utilize efficient, parallel space–channel context. All these codecs are limited to a single rate–distortion setting per checkpoint. In contrast, flexible-rate models are using approaches such as conditioning on the rate parameter [16], spatially adaptive quality maps [47, 53], and hierarchical VAEs with quantization-aware priors [19].
@@
Beyond images, learned video compression targets temporal redundancy across frames [1, 33, 34]. While classical approaches rely on motion estimation and compensation, transformer-based models remove such priors and model temporal dependencies directly in latent space. The Video Compression Transformer [39] follows this design, encoding frames independently and using a temporal transformer to predict latents from past context, making it better suited to settings with limited or irregular motion. A complementary line of research investigates Implicit Neural Representations (INR), which fit a small network to each image or video and store the signal in its weights, achieving strong rate–distortion performance but requiring per-sample optimization [30, 21, 63, 35].
@@
Earth observation data. Most neural compression targets natural imagery, whereas EO includes multispectral bands, higher bit depth, and long temporal horizons. Compression must preserve spectral and structural cues relevant for downstream analysis [23], aligning with the broader paradigm of task-oriented compression [54, 46]. Operational EO pipelines typically rely on the JPEG2000 and CCSDS standards [61, 12, 11] for their robustness and low complexity. Learned models have been explored for optical and SAR images [37, 18], with a focus on reducing on-board complexity [3] and spectral grouping [28]. Other works exploit spatial–spectral encoders and mixed hyperpriors to capture redundancy [31, 58, 20, 22]. Recent generative approaches focus on low-bitrate RGB imagery using diffusion models [60, 64], and INR–based methods have been explored for multispectral data [15].
Despite this progress, EO compression research has predominantly focused on single-image settings and mostly relies on RGB or small-scale datasets, with limited exploration of temporal modeling. To our knowledge, no pretrained models are publicly available for the widely used Sentinel–2 imagery. TerraCodec aims to fill these gaps and offers multispectral neural codecs, a temporal transformer model to capture long-range dependencies, and single-checkpoint, flexible-rate compression.
@@ 3 Methodology
We begin with an overview of our EO compression approach, then detail the architectures of the TerraCodec models (Sec. 3.1), and finally introduce Latent Repacking for flexible-rate models (Sec. 3.2).
We study lossy compression of multispectral, multi-temporal EO imagery. An EO sequence is a set of images $\mathbf{x}_{i}\in\mathbb{R}^{H\times W\times C}$, each of size $H\times W$ with $C$ spectral bands. While EO sensors range from a single panchromatic channel to finely sliced hyperspectral imagers, we focus on Sentinel–2 L2A with C=12C{=}12 optical bands in the visible and near infrared, saved as 16-bit reflectance.
@@
EO-specific choices. Our codecs adopt three EO-specific design choices: (i) native support for 12-band, 16-bit inputs; (ii) pretraining on a large-scale global EO dataset; and (iii) per-band standardization rather than global normalization, to stabilize training and preserve band-specific statistics for downstream tasks. Following prior literature [5, 39], we adopt CNN-based encoders for intra-frame compression due to their efficiency and low latency.
@@ 3.1 TerraCodec
We introduce TerraCodec, a family of learned codecs for EO, including two image codecs: a lightweight factorized prior (TEC-FP) and a stronger space–channel context model (TEC-ELIC), as well as a temporal transformer (TEC-TT), with a flexible-rate variant (FlexTEC).
Figure 2: TerraCodec image codecs: Factorized Prior [5] uses a fully factorized prior without context, and assumes zero-centered normal latents. ELIC [26] augments a hyperprior with spatial and channel context to predict per-latent mean/scale.
Factorized Prior (TEC-FP). TEC-FP is a Factorized Prior model [5], our most basic TEC image codec. It employs a fully factorized entropy model, where each element of the quantized latent $\hat{\mathbf{y}}$ is modeled independently by $q_{\phi}(\hat{y}_{j})$, without side information or context. It is illustrated in the upper part of Figure 2. This yields fast, parallel entropy coding, but limited expressiveness compared to hyperprior- or context-based models.
Efficient Learned Image Compression (TEC-ELIC). TEC-ELIC instantiates ELIC’s space–channel context entropy model [26] for EO inputs. The encoder-decoder network includes residual bottleneck and attention blocks, increasing representational capacity. The entropy model predicts per-latent mean and scale from (i) spatial context via checkerboard convolutions, (ii) channel context from previously decoded latent groups, and (iii) side information from a hyperprior, improving rate–distortion performance at the cost of higher complexity. Figure 2 illustrates the hyperprior context model in simplified form.
Figure 3: Architecture of the TerraCodec-$T$ model following VCT [39]. Each input image is first encoded by an ELIC encoder. The per-image latents are tokenized, and a temporal transformer models these tokens autoregressively, predicting the mean and scale parameters for the current frame based on past latents.
Temporal Transformer (TEC-TT). TEC-TT builds on the VCT architecture [39]. We train a transformer to model temporal dependencies of seasonal EO data in latent space, predicting the current frame’s latent distribution from past context. Each frame $\mathbf{x}_{i}$ is encoded to latents $\mathbf{y}_{i}$ and quantized. We partition the current latent into $B$ non-overlapping spatial blocks $\{\hat{\mathbf{y}}_{i,b}\}_{b=1}^{B}$ and the two past latents into overlapping context blocks to increase the transformer’s receptive field. Each block $b$ is flattened into a sequence of $T$ tokens $\{\hat{y}_{i,b,t}\}_{t=1}^{T}$ of channel width $d_{\text{lat}}$. A temporal encoder aggregates the two previous frames into a joint context embedding $z_{\mathrm{joint}}=E(\hat{\mathbf{y}}_{i-2},\hat{\mathbf{y}}_{i-1})$. As shown in Figure 3, within each current block, a masked autoregressive transformer predicts token-wise prior parameters conditioned on already decoded tokens and $z_{\mathrm{joint}}$:
@@
We assume conditional independence across blocks given the context, allowing parallel probability estimation during encoding and parallel block decoding. Causal masking prevents attention to undecoded tokens; see the supplementary material for details. TEC-TT uses the same CNN analysis–synthesis transforms as TEC-ELIC. Unlike the original VCT, it is trained end-to-end on the rate–distortion objective without image pretraining, using a $\lambda$-schedule that emphasizes low-rate regimes early. We further adapt TEC-TT for flexible-rate scaling, introducing the FlexTEC variant in the next section.
@@ 3.2 Latent Repacking for flexible-rate models
From spatial tokens to channel slices. A standard transformer codec represents an image block with $T$ spatial tokens, each spanning the full latent dimension $d_{\text{lat}}$. Dropping tokens at inference discards entire regions, causing severe artifacts (see the supplementary material). Instead, we aim for early tokens to encode information that is globally useful across the scene.
The latent block can be viewed as a 3D tensor $A\in\mathbb{R}^{H\times W\times d_{\text{lat}}},$ with $H\times W=T$ tokens and latent dimension $d_{\text{lat}}$. In the standard layout, each token $t$ is a spatial patch $(h,w)$ containing all $d_{\text{lat}}$ channels at that location. Latent repacking instead slices the channel axis into $T$ groups of width $k=\tfrac{d_{\text{lat}}}{T},$ and redefines tokens so that each spans the full scene but only $k$ channels.
@@
In other words, $t^{\prime}_{u}$ is the $u$-th slice of $k$ consecutive channels of $A$, spanning the full spatial field. The procedure is reversible; reapplying slicing and repacking restores the original layout. After repacking, keeping the first $K$ tokens corresponds to $A[:,:,0:K\cdot k],$ the first $K\cdot k$ latent channels at every spatial location.
@@
To keep later tokens informative, we use dynamic rate scaling: when fewer tokens are kept, their rate loss weight is upweighted. It prevents information from collapsing into the first tokens and encourages useful content across all tokens. Budgets $K$ are sampled more frequently at higher values during training (as in FlexTok [4]), ensuring that all tokens are trained while the model also learns to operate across a range of rates.
We apply the approach on TEC-TT and introduce the Flexible-Rate TerraCodec (FlexTEC) model. The model applies Latent Repacking and masking inside the temporal transformer after image-wise compression and restores the original layout before image decoding (details in the supplementary material).
@@ 4 Experimental setup
This section describes the data and pretraining (Sec. 4.1), evaluation and baselines (Sec. 4.2), as well as downstream tasks (Sec. 4.3) used to assess TerraCodec.
@@ 4.1 Pretraining
All TEC models are pretrained on SSL4EO-S12 v1.1 [9, 56], a large-scale Sentinel–2 corpus with 244k globally distributed locations and four seasonal snapshots per location. Each L2A sample consists of $\times264$ pixels at 10 m resolution; we crop to $\times256$ pixels (random crops for training, center crops for evaluation) to ensure uniform size. Bands at 20 m and 60 m are upsampled to 10 m with nearest-neighbor interpolation for spatial alignment. We follow the official spatial split into training and validation sets. To stabilize multispectral training, each input band $b$ is standardized by its dataset mean $\mu_{b}$ and standard deviation $\sigma_{b}$. Losses are computed in this standardized space, which balances gradient magnitudes across channels and avoids overfitting to high-variance bands.
Image codecs. TEC-FP and TEC-ELIC are optimized with Adam (learning rate $10^{-4}$), using an auxiliary learning rate of $5\cdot 10^{-3}$ for the entropy bottleneck, gradient clipping at 1.0, and mixed precision. We employ a cosine learning-rate schedule with 5% warmup and $\eta_{\min}=10^{-5}$. Models are trained for 100 epochs with batch size 64 on a single NVIDIA A100 GPU, requiring 20–25 hours. A temporal index is randomly sampled for each sample in every epoch, so that one epoch covers 25% of the dataset. We sweep five $\lambda$ values to span low- to high-bitrate regimes.
Temporal codec. TEC-TT is trained with a temporal context of two past frames for 300k steps with a global batch size of 24 on four NVIDIA A100 GPUs, requiring about 70 hours. We optimize with AdamW ($\text{lr}\!=\!10^{-4}$, weight decay $10^{-2}$) and employ half-cosine decay schedule to $\eta_{\min}=10^{-6}$ with 15% warmup steps. We sweep six $\lambda$ values for varying rate–distortion settings. For low-rate settings ($\lambda\leq 5.0$), we scale $\lambda$ by 10 during the first 15% of training before annealing back to the target value. Further pretraining details, including FlexTEC, are given in the supplementary material.